• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name M Jones

    Log entry time 16:09:01 on November14,2002

    Entry number 90099

    keyword=day shift summary

    Doing j50 kinematics, Roc 4 has been disabled in DAQ becuase of the fastbus
    crate died so no fpp data.
    Run 3166 was going when shift started.
    Run 3166 finished at 8:34 1M events
    Run 3167 started at 8:35 finished at 9:04 0.58M events
    Run 3168 started at 9:05 finished at 9:39 0.58M events

    Beam down for Beam Studies. Access, Ed Folts moves Big bite to correct angle
    and Riad puts a detector on the pivot area.

    Close up at 11:30 . No beam until 13:00
    Run 3168 started at 13:08 finished at M events
    Run 3169 (1.1 M events)
    Run 3170 (1 M events)
    Run 3171 (0.96M)

    15:55: Beam lost - MCC called and needs to reboot IOC.